OpenStudy (anonymous):

MEDAL! 3/4 divided by 5/18 Find the quotient and simplify your answer.

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

helpppppp

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 } \div \frac{ 5 }{ 18 } = \frac{ 3 }{ 4 } \times \frac{ 18 }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 } \times\frac{ 9 }{ 5 }\]

OpenStudy (anonymous):

Now multiply top and bottom.